Under the tax rate of 40%, an unlevered firm Pacific Company’s WACC is currently 15 percent. The company can borrow at 6 percent.

a.   What is Pacific Company’s cost of equity? 

b.   If the firm converts to 30 percent debt, what will its cost of equity be? 

c.   If the firm converts to 55 percent debt, what will its cost of equity be? 

d.   What is Pacific Company’s WACC in part (b)? In part (c)?   



1
Expert's answer
2022-02-28T11:42:22-0500

Question 1

WACC= "\frac{E}{V}\times K_e + \frac{D}{V}\times K_d(1-tax)"

Where E= Value of equity 

D=Value of debt

"K_e" = Cost of equity 

"K_d" = Cost of debt 

V= Total value of firm = Debt+Equity

By substituting given values in above formula we get,

(a) The firm is unlevered which means there is no debt in its capital structure so debt will be 0.

"0.14=1\times K_e"

"K_e" = 0.14 or 14%

 

(b)"K_e=K_o+(K_o-K_d)\frac{D}{E}\times(1-Tax)"

="0.14+(0.14-0.06)\frac{0.3}{0.7}(1-0.40)"

=0.14+0.0206=0.1606 or 16.06%

 

(c) D="55\%" so E=1-D="45\%"

"K_e" =0.14+(0.14-0.06)0.55/0.45(1-0.4)

=0.14+0.0586=0.1986 or 19.86%

 

(d) 

WACC when D=0.30

WACC="0.1606\times 0.70+0.30\times0.06\times(1-0.4)"

WACC=0.11242+0.0108=0.1232 or 12.32%

 

WACC in (c) Where D=0.55

WACC="0.1986\times0.45+0.55\times0.06\times(1-0.4)"

WACC=0.08937+0.0198=0.10917 or 10.92%